OpenStudy (anonymous):

(x^2)/(x-5) - 1 = (25)/(x-5) solve for x

OpenStudy (anonymous):

\[\frac{x^2}{x-5}-1=\frac{25}{x-5}\] right?

OpenStudy (anonymous):

I think the -1 is in the denominator of the first rational expression.

OpenStudy (anonymous):

its not. Satellite is correct

OpenStudy (anonymous):

i would subtract \(\frac{25}{x-5}\) from both sides and add 1 to both sides and start with \[\frac{x^2-25}{x-5}=1\]

OpenStudy (anonymous):

you can just about solve this in your head, but lets take the steps multiply both sides by \(x-5\) to get \[x^2-25=x-5\] factoring gives you \[(x+5)(x-5)=x-5\] divide and get \[x+5=1\] another method would be to note that \[\frac{x^2-25}{x-5}=\frac{(x+5)(x-5)}{x-5}=x+5\] so you had \[x+5=1\] all along now you can solve for \(x\) in one step

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Don't forget to close your questions please.

OpenStudy (anonymous):

I think there's one more answer. x^2/(x-5) - (x-5)/(x-5) = 25/(x-5). (x^2 - x + 5)/(x-5) = 25(x-5). Multiply both sides by (x-5). x^2 - x +5 = 25. x^2 - x -20 = 0. (x-5)(x+4) = 0. Therefore, x=5 is answer as well as x=-4.

OpenStudy (anonymous):

Ah, forgot that it makes us divide by 0, my bad.